# iPanda: An Intelligent Protocol Testing and Debugging Agent for Conformance Testing

### v1

@@ 4.1.4. Baseline
Specification-driven conformance testing methods typically require the development of protocol-specific conformance testing tools. Unfortunately, no open-source conformance testing tools currently exist for CoAP and RSocket. Moreover, our work is the first to employ LLMs for conformance testing in a protocol-agnostic manner. To reasonably evaluate iPanda’s performance, we selected a pure-LLM baseline approach in which the LLM generates test code in a single step (using the same prompt and randomness parameter settings). This corresponds to the startup stage of iPanda with the memory module disabled, serving as our experimental baseline.
4.1.5. Metrics
@@
. This metric is commonly used to evaluate the correctness and reliability in generating code for a given programming task, which is a core sub-task of iPanda. measures the probability that at least one of the generated code solutions is correct. Given the inherent randomness in LLM-generated outputs, we can simplify this metric. Specifically, for a test set containing cases , the metric for a LLM is defined as:
@@
where denotes the XOR operator, and represents the correctness evaluation of the -th generated solution. It takes a value of if the solution is correct and otherwise.
Conformance testing results. This is a qualitative metric that includes both positive sample (generating executable code) and negative sample (generating unexecutable code due to library bugs possibly) analyses based on whether the protocol implementation complies with the specification requirements. It is used to demonstrate the effectiveness of iPanda in conformance testing.
4.2. Marginal benefit analysis in the iteration stage
@@
As the number of iterations increases, the marginal benefit of code correction in the iteration stage gradually decreases. To achieve the optimal marginal benefit, we must first determine an appropriate maximum step of LLM iterative reasoning. In our experiments, we set the upper limit for reasoning iterations to . iPanda was tested on the CoAP-set, which contains a larger number of test samples. To analyze the required number of reasoning iterations for successfully generating executable code, we constructed a histogram depicting the cost of successful generations. The experimental results are shown in Fig. 5. It can be observed that in most cases, no more than steps are required to generate executable code. Specifically, when setting the maximum inference iterations to , using GPT-4o, iPanda successfully tested cases in the CoAP-set (classified as positive samples), among which cases required no more than iterations, accounting for . Using DeepSeek-V3, the corresponding proportion was . Using Qwen2.5-Coder-32B, the proportion was . Under the influence of the scaling law, large-scale foundation models such as GPT-4o and DeepSeek-V3 (both exceeding parameters) exhibit stable performance in code generation and correction. As the number of iterations increases, the marginal benefit steadily declines. However, for Qwen2.5-Coder-32B, which has only parameters, the marginal benefit is relatively unstable. This may be attributed to the model’s smaller scale, leading to weaker code generation capabilities. By default, we set the maximum step of LLM reasoning iterations to in subsequent experiments to achieve the optimal marginal benefit.
4.3. Performance analysis of code generation
@@
We conducted experiments using the baseline and iPanda supported by the three LLMs to evaluate their code generation performance on the CoAP-set. The experimental results for all methods are shown in Tab. 2. Experimental results show that under the same LLM conditions, iPanda significantly enhances code generation capabilities compared to the baseline. is improved by (GPT-4o), (Deepseek-V3), and (Qwen2.5-Coder-32B) respectively. For example, when using the GPT-4o model, the baseline achieves only a of . In contrast, iPanda improves this by , reaching . This strongly validates that aiocoap aligns with the specifications underlying these test cases, reducing the likelihood of nonconformances and narrowing the focus for critical testing. Even with the smallest model Qwen2.5-Coder-32B, iPanda’s is nearly twice that of the baseline using the most advanced model GPT-4o. These results confirm the success of the augmented CRITIC method. With its support, even smaller-scale LLMs can be stimulated to exhibit stronger code generation capabilities, achieving performance comparable to or even surpassing larger, more advanced LLMs.
@@
Due to the inherent randomness in LLM-generated content, multiple repeated tests are typically required for code generation, and performance is often evaluated using . Therefore, we conducted a joint experiment to examine the impact of the maximum reasoning iterations and the number of repeated tests . Given the strong performance of GPT-4o, we selected it as the default LLM for iPanda. The experiment was configured with and . The experimental results on the CoAP-set are shown in Fig. 6-(a), where the numbers in squares represent the number of executable code generated by iPanda under the current and . When setting and , the experiment degenerates into the setup of Fig. 5-(a). When setting and , the experiment degenerates into the baseline evaluation. From the results, we observe that increasing the reasoning iterations significantly improves iPanda’s performance. Additionally, repeated testing can also enhance the success rate of code generation, though its impact is considerably smaller than increasing reasoning iterations. For instance, increasing the number of repeated tests by five (i.e., instead of ) raises the number of positive samples from to . Increasing the number of inference iterations by five (i.e., instead of ) raises the number of positive samples from to . These results validate the rationality of our decision to adopt a sequential strategy (i.e., iterative code generation using a single LLM), rather than a parallel strategy (i.e., involving multiple LLMs generating code simultaneously). While the former has a higher interaction cost per step due to increased context length, this additional context contains valuable execution feedback, which provides richer information to guide the LLM in refining the generated code.
4.4. Ablation study
To validate the effectiveness of the methods used in iPanda, we conducted an ablation study to analyze the contributions of codeRAG and augmented CRITIC. The experimental results are presented in Tab. 3.
@@ 4.4. Ablation study
| iPanda, no codeRAG | 80.95% | 14.51% |
| iPanda, no augmented CRITIC | 17.32% | 3.23% |
@@
Since LLMs have a deep understanding of CoAP and aiocoap, the long-term memory repository was frozen, effectively disabling the codeRAG mechanism. As a result, codeRAG does not impact the testing outcomes for this protocol, leading to only two experimental configurations in the ablation study. To evaluate the effectiveness of codeRAG, we introduced additional tests for the RSocket protocol and its Python implementation, rsocket-py. In this setup, iPanda was preloaded with the open-source code of rsocket-py for use in codeRAG. Experimental results on the RSocket-set revealed that the inclusion of codeRAG significantly improved LLM’s ability to understand and correctly utilize rsocket-py. This resulted in an increase in from to .
@@
When the augmented CRITIC is removed, iPanda degrades into a baseline method that only incorporates codeRAG. Experimental results show that, on both datasets, the performance of iPanda drops significantly without the augmented CRITIC. This phenomenon has been extensively studied in previous experiments. Nevertheless, on the RSocket-set, even without the augmented CRITIC, iPanda still outperforms the baseline method. This further demonstrates the effectiveness of codeRAG.
@@
Remark: The primary design goal of iPanda is to support conformance testing for multiple communication protocols. Therefore, it must exhibit protocol compatibility. At the same time, we also expect iPanda’s performance to improve as the underlying LLMs advance, requiring it to maintain model compatibility. The extensive performance and ablation experiments conducted above confirm that iPanda demonstrates strong compatibility with both different protocols and different LLMs.
4.5. Results analysis of conformance testing
@@
According to the joint experiment results in Sec. 4.3, maximizing the number of iterations and the number of repetitions yields the highest number of positive samples, indicating that aiocoap conforms to the majority of protocol specifications. For negative samples, we analyzed the documents they belong to, as shown in Fig. 6-(b). Notably, Most of the negative samples are concentrated in the CoAP’s RFC 9177. This suggests that aiocoap has likely not yet implemented the specifications outlined in RFC 9177. This hypothesis aligns with the author’s introduction on GitHub (aiocoap, 2025), which specifies the standards supported by aiocoap, confirming that RFC 9177 has not yet been implemented in aiocoap. By generating as many positive samples as possible, iPanda can effectively reduce the scope of conformance testing. In addition, the test results of negative samples also guide further testing.
5. Related work
@@
Conformance testing tools. Several tools have been developed for communication protocol conformance testing. Testing and Test Control Notation version 3 (TTCN-3) is a widely used language for rigorous conformance certification in mobile communications, and IoT protocols (Grabowski et al., 2003). Scapy enables flexible construction and transmission of custom packets via scripting, aiding tests of protocol edge cases and exception handling, particularly in security assessments (Rohith et al., 2018). Protocol fuzz testing tools, such as Fairfuzz and Boofuzz, have demonstrated strong capabilities in uncovering vulnerabilities and testing robustness (Pereyda, 2019; Lemieux and Sen, 2018; Peng et al., 2018). However, these tools are typically applied at specific stages within existing testing workflows and heuristic-based approaches (Makhmudov et al., 2025). A major limitation remains their dependence on the manual creation of test cases and scripts, requiring substantial developer effort. Currently, no tool seamlessly integrates protocol documentation, implementation, and tests into an end-to-end conformance testing process. Addressing this gap is precisely the focus of our work.
@@
LLM-Based automated testing tools. Automated testing tools based on LLMs have gained increasing attention in recent years. Researchers have begun exploring LLM-powered agents for automating testing tasks. In code testing, some studies leveraged LLMs to generate high-coverage unit tests, such as those for the JUnit testing framework (Siddiq et al., 2024; Guilherme and Vincenzi, 2023). LLIFT (Li et al., 2023) employs LLMs for static code analysis to detect potential security vulnerabilities. PENTESTGPT, an LLM-based penetration testing tool, effectively identifies common vulnerabilities and analyzes source code for flaws (Deng et al., 2023b). DB-GPT integrates the Tree-of-Thought method into LLMs to systematically analyze database anomalies (Zhou et al., 2023). In fuzz testing, existing tools often struggle with message format obfuscation and dependencies between messages. To address this, LLMIF (Wang et al., 2024) integrates LLMs into IoT fuzz testing to automate the extraction of protocol formats and device response inference. However, current LLM-driven testing tools typically target isolated tasks and do not effectively support conformance testing. To address this research gap, we introduce iPanda, an LLM-based agent tailored specifically for protocol conformance testing.
@@
Augmented LLMs. Although LLMs excel in tasks such as question answering and text generation, their performance remains constrained by the limitations of their pre-trained datasets and the contextual information provided during inference. As a result, they may perform poorly in certain specialized tasks. To address these limitations, researchers have explored various tools to enhance LLM capabilities, including Web browsers (Deng et al., 2023a; Nakano et al., 2021; Chen et al., [n. d.]), RAG (Du et al., 2024; Jeong, 2023; Ng et al., 2025), programming tools (Alinezhad et al., 2019; Lu et al., 2023), other deep neural network models (Shen et al., 2023), and so on (Wang et al., 2022; Yao et al., 2023; Shinn et al., 2023). Similarly, in our work, iPanda incorporates RAG and programming tools to enhance the effectiveness of LLMs, improving their ability to analyze protocol specifications, generate test cases, and interact with protocol implementation libraries.
6. Conclusion
@@
We present iPanda, an LLM-based intelligent agent for automated, end-to-end conformance testing of communication protocols. iPanda leverages the keyword-based TCG method to efficiently generate test cases from protocol documents, and employs code-oriented prompting methods (e.g., codeRAG and Augmented Role Prompting) to enhance its understanding of tested protocol implementation and generate test code. Furthermore, the augmented CRITIC is integrated for iterative code refinement. Through execution-based validation, iPanda assesses whether the protocol implementation adheres to the specified protocol requirements. Experimental results prove the effectiveness and high efficiency of iPanda in conformance testing.

### v2

@@ IV-A4 Baseline
Specification-driven conformance testing methods typically require the development of protocol-specific conformance testing tools. Unfortunately, no open-source conformance testing tools currently exist for CoAP and RSocket. Moreover, our work is the first to employ LLMs for conformance testing in a protocol-agnostic manner. To reasonably evaluate iPanda’s performance, we selected a pure-LLM baseline approach in which the LLM generates test program in a single step (using the same prompt and setup). This corresponds to the startup stage of iPanda with the memory module disabled, serving as our experimental baseline.
IV-A5 Metrics
@@
Pass@k. This metric is commonly used to evaluate the correctness and reliability in generating code for a given programming task, which is a core sub-task of iPanda. measures the probability that at least one of the generated code solutions is correct. Specifically, for a test set containing cases , the metric for an LLM is defined as:
@@
where denotes the XOR operator, and represents the evaluation of the -th generated solution. It takes a value of if the solution is correct and otherwise.
Conformance testing results. To evaluate the effectiveness of iPanda in conformance testing, we conduct a qualitative assessment of the generated programs compliance with specifications. The assessment categorizes results into two types:
Positive samples: The generated program is executable and functions in accordance with the specification.
Negative samples: The code fails to execute or functionally violates the protocol requirements.
IV-B Marginal Benefit Analysis
@@
As the number of iterations increases, the marginal benefit of program correction gradually decreases. To achieve the optimal marginal benefit, we must first determine an appropriate maximum step of LLM iterative reasoning. In our experiments, we set the upper limit for reasoning iterations to . iPanda was tested on the CoAP-set, which contains more test samples. To analyze the required number of reasoning iterations for successfully generating executable program, we constructed a histogram depicting the cost of successful generations. The experimental results are shown in Fig. 3. It can be observed that in most cases, no more than steps are required to generate executable program. Specifically, when setting the maximum step to , using GPT-4o, iPanda successfully tested cases in the CoAP-set (classified as positive samples), among which cases required no more than iterations, accounting for . Using DeepSeek-V3, the corresponding proportion was . Using Qwen2.5-Coder-32B, the proportion was . Under the influence of the scaling law, large-scale foundation models such as GPT-4o and DeepSeek-V3 (both exceeding parameters) exhibit stable performance in program generation and correction. As the number of iterations increases, both marginal benefit steadily declines. However, for Qwen2.5-Coder-32B, which has only parameters, the marginal benefit is relatively unstable. This may be attributed to the model’s smaller scale, leading to weaker program generation capabilities. Resultly, we set the maximum step of iterations to in subsequent experiments to achieve the optimal marginal benefit.
IV-C Performance Analysis of Program Generation
@@
We conducted experiments using the baseline and iPanda supported by the three LLMs to evaluate their code generation performance on the CoAP-set. The experimental results for all methods are shown in Tab. II. Experimental results show that under the same LLM conditions, iPanda significantly enhances program generation capabilities compared to the baseline. is improved by (GPT-4o), (Deepseek-V3), and (Qwen2.5-Coder-32B) respectively. For example, when using the GPT-4o, the baseline achieves only a of . In contrast, iPanda improves this by , reaching . This strongly validates that aiocoap aligns with the specifications underlying these test cases, reducing the likelihood of nonconformances and narrowing the focus for critical testing. Even with the smallest model Qwen2.5-Coder-32B, iPanda’s is nearly twice that of the baseline using the most advanced model GPT-4o. These results confirm the success of the iterative optimization algorithm. With its support, even smaller-scale LLMs can be stimulated to exhibit stronger program generation capabilities, achieving performance comparable to or even surpassing larger, more advanced LLMs.
@@
Due to the inherent randomness in LLM-generated content, multiple repeated tests are typically required for program generation, and performance is often evaluated using . Therefore, we conducted a joint experiment to examine the impact of the maximum reasoning iteration step and the number of repeated tests . Given the strong performance of GPT-4o, we selected it as the default LLM for iPanda. The experiment was configured with and . The experimental results on the CoAP-set are shown in Fig. 4-(a), where the numbers in squares represent the number of executable program generated by iPanda under the current and . When setting and , the experiment degenerates into the setup of Fig. 3-(a). When setting and , the experiment degenerates into the baseline evaluation. From the results, we observe that increasing the reasoning iteration step significantly improves iPanda’s performance. Additionally, repeated testing can also enhance the success rate of program generation, though its impact is considerably smaller than increasing reasoning iterations. For instance, increasing the number of repeated tests by five (i.e., instead of ) raises the number of positive samples from to . Increasing the number of inference iterations by five (i.e., instead of ) raises the number of positive samples from to . These results validate the rationality of our decision to adopt a sequential strategy (i.e., iterative program generation using a single LLM), rather than a parallel strategy (i.e., involving multiple LLMs generating program simultaneously). While the former has a higher interaction cost per step due to increased context, this additional context contains valuable execution feedback, providing richer information to guide the LLM in refining the generated program.
IV-D Ablation Study
To validate the effectiveness of the methods used in iPanda, we conducted an ablation study to analyze the contributions of code-oriented RAG and iterative optimization algorithm. The experimental results are presented in Tab. III.
@@ IV-D Ablation Study
| iPanda, w/o code-oriented RAG | 80.95% | 14.51% |
| iPanda, w/o iterative optimization | 17.32% | 3.23% |
@@
Since LLMs have a deep understanding of CoAP and aiocoap, the long-term memory repository was frozen, disabling the code-oriented RAG. As a result, RAG does not impact the testing outcomes for this protocol, leading to only two experimental configurations in the ablation study. To evaluate the effectiveness of code-oriented RAG, we introduced additional tests for the RSocket protocol and its Python implementation, rsocket-py. In this setup, iPanda was preloaded with the open-source code of rsocket-py for use in RAG. Experimental results on the RSocket-set revealed that the inclusion of code-oriented RAG significantly improved LLM’s ability to understand and correctly utilize rsocket-py. This resulted in an increase in from to .
@@
When the iterative optimization algorithm is removed, iPanda degrades into the baseline that only incorporates code-oriented RAG. Experimental results show that, on both datasets, the performance of iPanda drops significantly without the iterative optimization algorithm. This phenomenon has been extensively studied in previous experiments. Nevertheless, on the RSocket-set, even without the iterative optimization algorithm, iPanda still outperforms the baseline. This further demonstrates the effectiveness of code-oriented RAG.
@@
Remark: The primary goal of iPanda is to support conformance testing for multiple communication protocols. Therefore, it must exhibit protocol compatibility. Meanwhile, we also expect iPanda’s performance to improve as the underlying LLMs advance, requiring it to maintain model compatibility. The extensive performance and ablation experiments conducted above confirm that iPanda demonstrates strong compatibility with both different protocols and different LLMs.
IV-E Results Analysis of Conformance Testing
@@
According to the joint experiment results in Sec. IV-C, maximizing the number of iterations and repetitions yields the highest number of positive samples, indicating that aiocoap conforms to the majority of protocol specifications. For negative samples, we analyzed the documents they belong to, as shown in Fig. 4-(b). Notably, Most of the negative samples are concentrated in the CoAP’s RFC 9177. This suggests that aiocoap has likely not yet implemented the specifications outlined in RFC 9177. This hypothesis aligns with the author’s introduction on GitHub [22], which specifies the standards supported by aiocoap, confirming that RFC 9177 has not yet been implemented in aiocoap. By generating as many positive samples as possible, iPanda can effectively reduce the scope of conformance testing. In addition, the test results of negative samples also guide further testing.
V Related work
@@
Traditional testing tools. Several tools have been developed for protocol testing. Testing and Test Control Notation version 3 (TTCN-3) is a widely used language for rigorous conformance certification in mobile communications, and IoT protocols [23]. Scapy enables flexible construction and transmission of custom packets via scripting, aiding tests of protocol edge cases and exception handling, particularly in security assessments [24]. Protocol fuzz testing tools, such as Fairfuzz and Boofuzz, have demonstrated strong capabilities in uncovering vulnerabilities and testing robustness [25, 26, 27]. However, these tools are typically applied at specific stages within existing testing workflows and heuristic-based approaches [28]. A major limitation remains their dependence on the manual creation of test cases and scripts, requiring substantial developer effort. Currently, no tool seamlessly integrates protocol documentation, implementation, and tests into one conformance testing process. Addressing this gap is precisely the focus of our work.
@@
LLM-based automated testing tools. Automated testing tools based on LLMs have gained increasing attention in recent years. Researchers have begun exploring LLM-powered agents for automating testing tasks. In code testing, some studies leveraged LLMs to generate high-coverage unit tests, such as those for the JUnit testing framework [29, 30]. PENTESTGPT, an penetration testing tool, effectively identifies common vulnerabilities and analyzes source code for flaws [31]. DB-GPT integrates the Tree-of-Thought method into LLMs to systematically analyze database anomalies [32]. In fuzz testing, existing tools often struggle with message format obfuscation and dependencies between messages. To address this, LLMIF [1] integrates LLMs into IoT fuzz testing to automate the extraction of protocol formats and device response inference. However, those LLM-driven testing tools typically target isolated tasks and do not effectively support conformance testing. RFCScan is most similar to our work. It builds semantic indexes for both implementation code and specifications to detect conformances between the code and the RFC specifications at the semantic level. However, this approach focuses on static comparisons between code and specifications, rather than simulating how developers write test cases to verify the actual runtime behavior of the protocol, which is precisely where our work differs. To address this research gap, we introduce iPanda, an LLM-based agent tailored specifically for protocol conformance testing.
@@
Augmented LLMs. Although LLMs excel in tasks such as question answering and text generation, their performance remains constrained by the limitations of pre-trained datasets and the context provided during inference. As a result, they may perform poorly in certain specialized tasks. To address these limitations, researchers have explored various tools to enhance LLM capabilities, including Web browsers [33, 34, 35], RAG [36, 37, 38], programming tools [20, 39], other deep neural network models [40], and so on [41, 42, 43]. Similarly, in our work, iPanda incorporates RAG and programming tools to enhance the effectiveness of LLMs, improving their ability to analyze protocol specifications, generate test cases, and interact with protocol implementation libraries.
VI Conclusion
@@
We present iPanda, the first LLM-based intelligent agent for automated conformance testing of communication protocols. iPanda leverages the keyword-based TCG method to efficiently generate test cases from protocol documents, and employs code-oriented RAG and customized CoT strategy to enhance its understanding of tested protocol implementation and generate test program. Furthermore, the ierative optimization algorithm is integrated for iterative program refinement in automated testing platform. Through execution-based validation, iPanda assesses whether the protocol implementation adheres to the specified protocol requirements. Experimental results prove the effectiveness and high efficiency of iPanda in conformance testing.